Privacy and Security Notice

Jefferson Lab
JLab Web Server 
Jefferson Lab Home Search Contact JLab

FEL Upgrade Project October 2002 Monthly Report

We have had a full month of installation and commissioning activities on the FEL upgrade.  Highlights include: (1) the linac is under vacuum from the gun chamber to last cryomodule;(2) installation and successful powering of the injector dipole magnets; (3) completion and sign-off of the operating procedure to allow us to power the SRF modules; (4) the Personnel Safety System (PSS) was certified to allow SRF commissioning of the new cryomodule; (5) completion and installation of the photogun light box assembly; (6) receipt of the ultrahigh vacuum chambers for the optical cavity mirror assemblies for both the IR and UV Upgrades; (7) the Gun High Voltage Power Supply was tested to full specification (10.1 mA) and (8) successful re-coating of the photogun ball electrode with the field emission suppression coating and successful dry assembly of the entire electrode structure, (9) the viewer assemblies and one of the optical diagnostics were installed in the wiggler and the wiggler vacuum chamber is being purged.

Jefferson Lab hosted the 2002 Industrial Physics Forum for the American Institute of Physics at Williamsburg on Oct. 27-29th with 170 attendees from industry, academia and government.  This meeting highlighted speakers on Monday Oct. 28th from recent users of the FEL during the 2000-2001 user run, and on Tuesday, Oct. 29th a session on “Science and Security” included the President’s Science Advisor, John Marburger, and the Director of Science and Technology from the Office of Naval Research, Jane Alexander.  The full program for the meeting is given at

Presentations were prepared for the Fall DEPS Symposium at Monterey, CA, on Nov. 12-14th. Two talks were sent in to the Navy for review: an overview of the JLab FEL program and a summary of FEL power scaling.

We hosted visit to JLab by Gil Graff and John Albertine from ONR on Thursday, Oct. 31 to discuss the FY02 cost and schedule performance and FY03 planning for the FEL project.

G. Neil was interviewed by the Science Office of the EPSRC of the UK as part of the Gateway review process of the 4GLS Project at Daresbury Laboratory.  They were presenting the "Business Case" to go get permission to proceed with full design and component development (similar to CD-1 in DOE-speak).

Project Cost Performance:
The Phase 2 IR Demo FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k.

Work for Phase 2 proceeded to plan during this reporting period.  Actual charges of $71k were accrued for October for total accrued costs of $4,305k to date for Phase 2.  At present the cost variance for IR Demo Phase 2 is $31k and the schedule variance is
- $137k.  We have met the planned cost at completion of Phase 2 effort.  All but 3% of the scheduled work (earned value) has been booked at the completion date.   Remaining items in the original project scope have not impeded the start-up of the FEL Upgrade commissioning which began in September. We are actively working with the vendors who are supplying the remaining hardware so that the delivery and integration of this hardware can be interleaved with the commissioning activities over the next 3 months and not impede the commissioning schedule.

The UV FEL project (Phase 1) was approved for $2,836k for a performance period from September 2001 to September 30, 2002.  Actual charges of $0k were accrued through October for a total accrued cost of $2,221k since project start.  Work performed for October was $32k for a total of $2,600k for the UV project to date, which results in 100% scheduled vs. 92% performed for the UV project.  At present the cost variance for the UV project (Phase 1) is $378k and the schedule variance is - $236k.  The current cost and schedule variances will clear when the last quarter of charges for the Aerospace subcontract are costed in December due to the 3 month late start on subcontract.

Phase 2 of the UV FEL Project was approved for  $1,472k for a performance period from August 2, 2002 to June 30, 2003.  The start date was 4 months later than planned due to late arrival of AF funds.  Actual charges of $153k were accrued through October for a total accrued cost of $1,231k since project start.  Work scheduled for October was $8k for a total of $24k for the UV (Phase 2) project to date, work performed for October was $393k for a total of $1,072k for the UV (Phase 2) project to date, which results in 2% scheduled vs. 73% performed for the UV (Phase 2) project.  At present the cost variance for the UV (Phase 2) project is -$159k and the schedule variance is $1,048k.  Because of the late arrival of funds significant expenditures had to made early in the first 2 months which have generated both the current cost and schedule variances.  Funds for this project are essentially fully obligated and we expect to track remaining earned value on the UV optical cavities and the Phase 2 Aerospace subcontract which is scheduled to be costed in the next quarter.

WBS 4 (Injector):
The gun assembly with all associated components, pumps and diagnostics was successfully assembled and pumped down.  The solenoid with correctors nested inside was mounted on the gun’s downstream vacuum tube.  Assembly of the light box was completed and it was installed onto its stand in the vault.  The ball cathode was successfully implanted and the gun electrode structure was successfully dry assembled.

Gun HVPS - Additional parts were designed, fabricated, and installed to ensure a good contact through the Running/Conditioning Resistor mechanism.  Resistors were added to the Glassman HVPS to enable operation to 600 kV.  The unit was tested and achieved 10.1 ma.  The FEL system specification calls for 10 ma from the Gun.

WBS 5 (SRF):
By the end of October, we had all the 7-cells cavity assemblies for the 3rd cryomodule welded and three of them tuned.  End group welding in progress and we are working hard to facilitate part flow through the machine shops to keep up with the e-beam welding schedule.

WBS 6 (RF):
A TOSP was written to allow the start of SRF commissioning without having the concrete blocks back in place at the roll-up door for a radiation shield.  Portable lead sheets, a radiation monitor connected to the PSS, and low power pulsed beam will be used to mitigate the hazard.

Zone 4 - The attachment of the HOM filters was completed this month.
All interlocks have been checked.   SRF Commissioning commenced.

Zone 2 - This zone is ready for SRF testing also.

WBS 8 (Instrumentation):
BLMs (Beam Loss Monitors) - Testing continues as the VME card firmware and EPICS applications are refined.  Everything looks good.  A picture of the 1st installed card set can be seen here:

Drive Laser Controls - The punch list for drive laser controls is mostly completed.  The RS232 cabling and controls testing has been done (Internal Shutter control and readback, water pump control, etc.).  Also, the drive laser controller's key switch status and water pump status have been fixed.  For the upgrade, a multifunction VME module (AcroMag card) has been implemented into iocfel10 (in the clean room) to handle these functions.  This was our first implementation of this product (which will also be used in other places/systems) and so the task included our learning curve as well.  With the Acromag card and the serial communications re-established, the drive laser is now ready for remote operation.

Ten (10) Beamviewer interface boxes were fabricated to support the rework of the old diagnostics to the new format and will be installed in the 0F and 1F regions.

Software progress: The beam viewer controls have been complete (for months), the magnet screens are being worked, the injector is complete as well as the injector power supply (these are running on upgraded serial communications to eliminate CAMAC), RF software is ready for commissioning, vacuum screens and controls are complete.  Work continues on the Linux IOC for video acquisition, and the DSP based on line FEL spectrometer.

All the components for the magnet communication upgrade have arrived and the install is complete.  Testing of the communications to the trim racks and the 5kW PS in the 0F region begin this month.  Hardware upgrades of the trim magnets racks in FL05B09and B10 have been completed.  Utility chassis modifications for magnet racks FL05B10 through13 have been completed and installed. Installation of the trim magnets on the beam line has begun in the 0F region.  Progress was made on the Injection Dipole 55 kW PS.  Testing of communications, hardware and software for the magnet trim rack in the 0F region (FL05B13) proved successful.  Trim magnets MGJ0F04 and MQJ0F05 were powered and current ramped on these magnets.

The new gun HV monitor and interlock chassis is complete and testing is underway. (F0117), the 8-channel fiber driver (F0118), the 16-channel isolation/relay I/O module

Fast Shutdown Down interlock and cabling for RF in zone 4 was installed this month.  Supporting software for FSD is complete and checkout is progress.  The 15 pin fanout boards for the 64-bit Interface Module have been received from the vendor and being populated in house.

Modifications to the Dump Skid junction box have been completed to support the new Conductivity Analyzers from Solu Comp.

We hosted a visit from Dr. Steve Smith of SLAC.  Steve is an expert on Beam Position Monitoring systems.  He advised us on techniques for the multipass BPM electronics that see both the accelerated and decelerated beams.  We also had lengthy discussions on the 8 button BPMs as well as the regular strip lines.

The Machine Protection System (MPS) is progressing well.  The initial programming is complete for the Beam Mode and Machine Mode input boards.  Front panels are being manufactured.  The fiber input (Fast Shut Down FSD) & output (to DLPC) board is in the final stage of layout.  .

The skew quad cabling from the 0F to the 3F regions have been routed and terminated.  Control screens for corrector magnets in the 1F and 2F regions and the Extraction dipole have been completed and waiting for checkout.  Field measurements were made on solenoid magnets MFF0F01 and MFF0F02 to test for polarity.  We spent some time to work on the magnet measurement test stand.  There are a number of cabling issues that have not been properly addressed for more than 10 years.  A number of the resolver and signal cables were re-terminated which solved the bulk of the intermittent problems that has been slowing down the throughput.

The EECad group has recently provided the majority of our backlogged circuit board projects.  Thanks Very Much!!!  As a result the check-prints and notes to revise and archive are beginning to pile up.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
The Purcell Gap sheet pieces are starting to be attached to the tops of the six magnets after the field clamps were pinned and the tooling ball holes drilled.
• The comparison test stand problem was solved by Kevin Jordan and his crew by changing out some cables that had bad terminations.
• An additional GW was placed on the stepper stand and started field clamp configuration for final measurements.  The measurement technicians have to learn new techniques to adjust these large field clamps.  After some preliminary adjustment tries, they ordered some simple custom tools made through the machine shop to make the adjustment process efficient.
Injector Dipoles (DU/DV)
• We again machined the pole face ends on the GUs because the first machining pass brought us just outside of the range that allows field clamp adjustment to bring the magnet to specified field integral.  After a final measurement, the magnet was released, installed and aligned along with the companion GVs in the extraction line.  This completes the magnet installation in the Linac for first commissioning.
• The last GU was measured and found to have an anomaly in the core field through the central cross section It has a warp to a portion of the Purcell gaps of the upper pole tip.  We will probably have to strip these sheets off and re glue them.  These were measured and passed by the vendor (Wang NMR) and so failed since we received them, perhaps during the edge machining.
Arc 180-Degree Dipoles (GY)
• Wang NMR has stepped up their work on coils to produce one GY by the end of November after getting the GQs and GXs ready for shipment.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• The first two units of each style of magnet are ready for shipment, awaiting a shipper’s truck.
• Work continued in potting the remaining 4 GX coils for the next shipment.
Trim Quad (QT)
• No additional progress to report.
Sextupole (SF)
• Master Machine continues making the shim parts.  They will receive the original SCs next Tuesday for disassembly.
Octupole (OT)
• We continue working on a manufacturing package.
Beam Line and Vacuum
• Master Machine continues in production status where one arc chamber a week is coming off the production line
• Injection line had all magnets after the quarter Cryomodule installed
• Chicane chambers continue fit up, welding and machining

WBS 11 (Optics):
We received the optical cavity vacuum vessels for the IR and UV Upgrades.  They are being uncrated, and will be worked on, one at a time, in our clean enclosure in User Lab 2.  The Survey and Alignment group has established the mounting coordinates for the optical cavities and their associated pedestals and girders for the alignment diagnostics.  The anchors are drilled, and the pedestals installed and aligned.  The OCMMS hardware is ready for installation and one beamline assembly was installed.  Hardware to establish a clean environment around the optical cavity vessels was received and will be installed once alignment activities in the area are complete.  The trolley hoists were received - once installed they will be used to lift the vacuum vessel "top hats" on and off the lower section.  This will make assembly and testing much easier.  The optical transport turning mirrors (turning cassettes 1, 3, & 4) are being checked and detailed.  The holder for the 3" outcouplers was signed and being procured.  This is also true for the insertable mirror.  All parts for the 0F region THz diagnostic were received.  They were test assembled then sent to be UHV cleaned.  They will be installed in the X-chamber at the earliest opportunity.

The outcouplers for first light, and high power operation at ~ 10 microns arrived from the vendor ~ 1 week ahead of schedule.  They are being inspected.  Assuming they pass inspection, we will have the optics required for first light.  The ultraviewers that were assembled were aligned and are ready for integration onto the support girder.  Installation of the first alignment HeNe for the optical hardware (and later, the wiggler and beamline viewers) began.

More test runs of the drive laser were made to test the EPICS controls.  A big thanks goes to the I&C group in their progress in resurrecting the software.  Since the optical transport isn't needed at this time we've left the vacuum pump connected to it.  We conducted another test run of the drive laser to test the EPICS controls.

We held a Commissioning Planning Meeting to establish operations ground rules as we bring up the hardware.  A number of issues were dealt with; initial decisions were made to assign one CEBAF operator to the team as an interface to future operation under CEBAF Operations control.  It was deemed inefficient to have all commissioning carried out by the CEBAF operators since system owners will have to characterize their own systems before they can be handed over for routine operation.  Establishment of configuration control was established as a high priority for each system once handed over.

A test procedure (TSOP) was written and approved to allow us to do SRF commissioning on off shifts while we continue installation and non-powered check-out activities during the first shift.  We thank the installation crew for preparing new portable lead shields for this activity. SRF commissioning began this month.

Aerospace/Jefferson Lab Subcontract
(for the Laser Microengineering Experimental Station )
We have initiated a procurement effort for development of the Engineering Model.  The major components of the system are as follows.  The optical tables and the active vibration isolation systems have been procured with Newport Corporation chosen as the vendor.  We have resolved that Aerotech Corp. of PA meets all the specifications required for our XYZ positioning systems.  Aerotech Corp. will also supply the motion control software.  We have chosen Spectra Physics Lasers as the vendor for our laser system.  Finally, Con-Optics Corp. has been chosen to be the vendor for the fast shutter pockels cells.  We expect these critical item units to arrive in the next 6-12 weeks.

Other Related Projects

IR Upgrade Commissioning (ONR)
This project was funded in FY02 with $1,000k, with a period of performance from July 1, 2002 to March 30, 2003.  Additional funding of 1,000k is anticipated in FY03.  In this reporting period, $275k has been expended for a total expenditure of $1,071k.  Commitments total $224k.
See Highlights section and WBS 4,6,8,9,10 and 11 reports above.

Cryomodule 3 (ONR)
This project was funded in FY02 with $400k.  Additional funding of $350k is anticipated in FY03.  In this reporting period, $68k has been expended for a total expenditure of $338k.  Commitments total $40k.
Technical:  See WBS 5 report above.

One micron wiggler project (ONR)
This project was funded in FY02 with $600k.  Completion funding of $600k is anticipated in FY03.  In this reporting period, $10k has been expended for a total expenditure of $130k.  Commitments total $9k.
We analyzed the use of the 1 micron wiggler for UV operation.  The conclusion is that it is possible but not cost effective.  We also analyzed using the existing optical klystron to produce 10 kW in the 1-2.2 micron wavelength range.  This looks quite reasonable for 1.6 and 2.2 micron operation and possible, though there is more risk for 1 micron operation.  We are in the process of preparing a tech note on this analysis because of the potential cost savings of using the existing wiggler for short wavelength operation.

JTO Cryomirror Project
During this reporting period expenditures of $35k were incurred.  Commitments total $121k.  Total expenditures to date are $263k.  The program is on track financially and is expected to deliver all elements of the SOW within projected cost. It is expected that this budget will be fully expended in December with Phase 1 elements complete.  We are prepared to continue on schedule with Phase 2 of the program.
We continued efforts for proceeding with engineering tests of the proposed system.
We ran into difficulties with the molybdenum holder that the sapphire mirror is brazed into. The holder had a machined "lip" to help capture the copper cooling tube during its high temperature braze step.  We discovered, after trying to have the Machine Shop bend the tube, and then our trying it ourselves, that the lip requires you "pry out " the loop far enough that the copper yields, causing a gap.  We decided other, external, fixturing would perform the same function, and removed the lip.  Two holders were modified in the way, and we are proceeding to have them plated and brazed.

We received two isolating holders, one of G10 glass epoxy, the other of Macor ceramic, to use for our cooling tests.  The G10, which is less brittle than the Macor, will be used in our first tests, where we may unduly stress it.  We will then switch to Macor, which is UHV compatible.

A sapphire flat was purchased and inspected for flatness. It will be edge-coated, then low-temperature brazed into the Mo holder. We will darken the central area so it will absorb a known amount of laser power when we test it at cryogenic temperatures.
Some literature searches were done to check on the value of dn/dT at cryogenic temperatures.  This has been measured, in part because a great deal of work on cryogenic TI-sapphire lasers requires it.  The literature search indicates that work is occurring for optics for LIGO-style interferometers, and by NASA for the next generation space telescope.  This work confirmed our belief that it is possible to fabricate the mirror surface at room temperature in such a way that it will have the correct figure at cryogenic temperatures.
Birefringence measurements were delayed while we worked the problems with the holder.  They will be conducted in November.

JTO Drive Laser Project
During this reporting period expenditures of $0.4k were incurred.  Commitments total $2.2k.  Total expenditures to date are $26k.  The program is on track financially and is expected to deliver all elements of the SOW within projected cost. It is expected that this budget will be fully expended in December with Phase 1 elements complete.  We are prepared to continue on schedule with Phase 2 of the program.
We are in the process of proceeding on the major laser hardware purchase.
We spent the month working through some technical details, among them, the timing jitter stabilization system used by our primary candidate, Time-Bandwidth.  We also spent some time working with our Procurement staff on the language in the sole source statement.  That, and work on the upgrade FEL optical cavity, prevented us from beginning the Procurement.  This was deferred to November.